I need help with solving a linear system of equation by the elmination method.
Solve the system of equations using addition/subtraction with multiplication method.
9x+5y=2
-11x+2y=30
The answer they give is (-2,4) but I don't get how they got it.
Can you please explain how to do it? Step by step? Please? Thanks

Answer by Maths68(1474)

9x+5y=2..............(1)
-11x+2y=30 ..........(2)
Multiply (1) by -2 and (2) by 5 then add both equations
-18x-10y=-4
-55x+10y=150
--------------------
-73x=146
Divide by -73 both sides
-73x/-73=146/-73
X=-2
Now put the value of x in any two equations (1) or (2)
9x+5y=2
9(-2)+5y=2
-18+5y=2
Add 18 both sides of the above equation
-18+18+5y=2+18
5y=20
Divide by 5 both sides
5y/5=20/5
y=4
(x,y)=(-2,4)
Check:
Put the value of x and y in equation (1) and (2)
9x9x+5y=2
9(-2)+5(4)=2
-18+20=2
2=2
-11x+2y=30
-11(-2)+2(4)=30
22+8=30
30=30